# Xuezhang Chen  and Shihong Zhang

### v1

@@ 1 Introduction
Suppose $n\neq 3$ and $\bar{f}_{i}\in C^{\infty}(\mathbb{S}^{n})$ for $i\in\{0,1,2,3\}$. For $\xi\in\mathbb{B}^{n+1}$ and $\eta\in\mathbb{S}^{n}$, we let
Then, for $i<j\in\{0,1,2,3\}$ we have
  • (i)
    For $i+j\neq 3$, the solution of the follow biharmonic boundary value problem
    $\begin{cases}\displaystyle\Delta^{2}U=0\qquad\quad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}U=\bar{f}_{i}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}U=\bar{f}_{j}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$ (1.7)
    is unique and has the form of
  • (ii)
    For pairs $(0,3)$ and $(1,2)$, problem (1.7) has multiple smooth solutions.
Remark 1.2.
It is notable that restricting $\xi$ to $\mathbb{S}^{n}$, $2\bar{P}_{3}^{3}(\xi,\eta)$ coincides with the Green function of fractional GJMS operator $P_{3}$ on $\mathbb{S}^{n}$ (cf. the first author and Shi [10, Theorem 1(3)]). As an application, we can recover a special case in Case [7, Theorem 1.4] about the relationship between extrinsic and intrinsic GJMS operators, see Proposition 3.1 and its biharmonic Poisson kernel proof in Appendix A. Moreover, the non-uniqueness in Theorem 1.2(ii) actually arises from the invertibility of fractional GJMS operators $P_{3}$ and $P_{1}$, see [7, 10] for the definition.
We are ready to introduce a biharmonic Green function on $\mathbb{B}^{n+1}$ with each of four proper pairs of conformal boundary operators. Let
$\Gamma(\xi-\eta)=\Gamma(|\xi-\eta|)=\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|\xi-\eta|^{3-n}\qquad\mathrm{~{}~{}when~{}~{}}n\neq 3$
and
$\Gamma(\xi-\eta)=\Gamma(|\xi-\eta|)=-\frac{1}{4|\mathbb{S}^{3}|}\log|\xi-\eta|\qquad\qquad\mathrm{~{}~{}when~{}~{}}n=3$
denote the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$. We consider an nonhomogeneous biharmonic boundary value problem with vanishing boundary conditions
$\begin{cases}\Delta^{2}U=\bar{f}\qquad&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \mathscr{B}_{i}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\\ \mathscr{B}_{j}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n}.\end{cases}$ (1.8)
Remark 1.3.
A clever observation in (4.1) provides a traditional way in understanding biharmonic Poisson kernel: Conformal boundary operators $\mathscr{B}_{i}^{3}$ indeed act as a bridge between the biharmonic Poisson kernel and fundamental solution.
Definition 1.2 (biharmonic Green function).
Let $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ and $i<j\in\{0,1,2,3\}$ with $i+j\neq 3$. A function $\bar{G}^{(i,j)}:(\xi,\eta)\in\overline{\mathbb{B}^{n+1}}\times\overline{\mathbb{B}^{n+1}}\mapsto\mathbb{R}\cup\{\infty\}$ for (1.8) is called a Green function, if $\bar{G}^{(i,j)}$ satisfies
We first notice that $\bar{G}^{(0,1)}$ coincides with the Boggio formula (see, e.g., Boggio [2] or Gazzola-Grunau-Sweers [15, Lemma 2.27]), next combine this with biharmonic Poisson kernel to derive explicit formulae of Green functions.
@@
For $n\neq 3$, the representation formula of Green function $\bar{G}^{(i,j)}$ to (1.8) is
where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{n+1}}$. Moreover, there holds
$0\leq\bar{G}^{(0,1)}(\xi,\eta)\leq\bar{G}^{(0,2)}(\xi,\eta)\leq\bar{G}^{(1,3)}(\xi,\eta)\leq\bar{G}^{(2,3)}(\xi,\eta),\quad\forall~{}\xi\neq\eta\in\overline{\mathbb{B}^{n+1}},$
In the critical dimension that $n=3$, we have
Theorem 1.4.
Suppose $i,j\in\{0,1,2,3\}$ and $i<j$. The following are explicit formulae of biharmonic Poisson kernels on $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
  • (1)
    Under the same assumptions as in Theorem 1.1 when $n=3$, then the same formulae of $P_{i}^{3}$ are true except for
    $P_{3}^{3}(X)=-\frac{1}{4|\mathbb{S}^{3}|}\log(t^{2}+|x|^{2})+C,\qquad C\in\mathbb{R}$
    for $X\in\mathbb{R}_{+}^{4}$, and the biharmonic Poisson kernel for (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$ when $i+j\neq 3$.
  • (2)
    Under the same assumptions as in Theorem 1.2 when $n=3$, then the same formulae of $\bar{P}_{i}^{3}$ are true except for
    for $\xi\in\mathbb{B}^{4},\eta\in\mathbb{S}^{3}$. Moreover, we have
    (i) $i=1,2$ and $j=3$, the solution of (1.7) is
    (ii) $(i,j)=(0,1)\mathrm{~{}or~{}}(0,2)$, the unique solution of (1.7) is
Theorem 1.5.
The representation formula of Green function $\bar{G}^{(i,j)}$ to (1.8) is
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right)+C,\quad C\in\mathbb{R};$
$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+C,\quad C\in\mathbb{R};$
for $\xi,\eta\in\overline{\mathbb{B}^{4}}$, where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{4}}$.
We shall assemble the biharmonic Poisson kernel part in Section 3 and the biharmonic Green function part in Section 4.
We next turn to prove classification theorems of nonnegative solutions to (1.5) in $\mathbb{R}_{+}^{n+1}$. Before presenting the proof of our results, in Section 2 we first need to study the nontrivial solutions of an associated homogeneous problem to (1.6) under suitable assumptions on the decay rate at infinity of the initial data, and also describe asymptotic behavior of two typical singular integrals originating from the biharmonic Poisson integrals for $\mathbb{R}_{+}^{n+1}$.
For simplicity, we introduce a ‘geometric bubble function’ associated to above biharmonic boundary value problem by
$U_{\varepsilon,x_{0}}(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\qquad(x,t)\in\mathbb{R}_{+}^{n+1},$
and its boundary function by $u_{\varepsilon,x_{0}}(x)=U_{\varepsilon,x_{0}}(x,0)$, where $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$.
For the pair $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, under appropriate decay rate at infinity of initial data, the intrinsic nature of $\mathscr{B}^{3}_{3}$ boundary condition enables us to determine the explicit representation of $u$ on $\partial\mathbb{R}_{+}^{n+1}$, see Lemma 5.1. Using basic properties of $U_{\varepsilon,x_{0}}$ and biharmonic Poisson kernel, we establish that every nonnegative smooth solution is a sum of a bubble function and a polynomial. When $p_{3}=p_{3}^{\ast}$, a striking feature is that the bubble function (in general, without geometric meaning) does not vanish even if $p_{1}\neq p_{1}^{\ast}$, and is exact $U_{\varepsilon,x_{0}}$ when $p_{1}=p_{1}^{\ast}$; whereas for $p_{1}<p_{1}^{\ast}$, the bubble function vanishes.
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{1}^{3}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some constant $c>\max\{3/p_{3},1/p_{1}\}$. Then we have
  • (1)
    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{1}>\frac{1}{n-3}$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{2}$ such that
@@
    $M_{1}(x,t)=\left[\mathbb{T}_{1}^{3}P_{1}^{3}\ast u_{\varepsilon,x_{0}}^{p_{1}}+\mathbb{T}_{3}^{3}P_{3}^{3}\ast u_{\varepsilon,x_{0}}^{p_{3}^{\ast}}\right](x,t).$
    In particular, if $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{2}t^{2}.$
  • (2)
@@
For a nonlinear boundary conditions involving the exceptional pair $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$, there still exists a surprising classification theorem for problem (1.5) relying not on $p_{1},p_{2}$ but on $p_{2}/p_{1}$.
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{1}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some constant $c>2/p_{2}$. Then we have
  • (1)
    If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$ and $p_{1}>0$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{3}$ such that
@@
    $M_{2}(x,t)=\left[P_{0}^{3}\ast u_{\varepsilon,x_{0}}^{\frac{p_{1}^{\ast}}{p_{1}}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast u_{\varepsilon,x_{0}}^{p_{2}^{\ast}}\right](x,t).$
    In particular, $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{3}t^{3}.$
  • (2)
@@
The pair $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$ becomes more challenging and $p_{3}$ is still the dominant exponent. Especially, when $p_{i}=p_{i}^{\ast}$ for $i=1,2,3$, the bubble function happens to be the same $U_{\varepsilon,x_{0}}$; meanwhile, the polynomial part varies according to distinct circumstances.
@@
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|X|\to\infty$ for some constant $c>\max\{2/p_{2},3/p_{3}\}$. Then we have
  • (1)
    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{1}$ such that
@@
    $M_{3}(x,t)=\left[P_{0}^{3}\ast u_{\varepsilon,x_{0}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast u_{\varepsilon,x_{0}}^{p_{2}}\right](x,t).$
    In particular, $p_{2}=p_{2}^{\ast}=\frac{n+1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{1}t.$
  • (2)
@@
These classification results demonstrate that the dominant impact of $\mathscr{B}^{3}_{i}$ is different, accordingly, $\mathscr{B}^{3}_{1}\preceq\mathscr{B}^{3}_{2}\preceq\mathscr{B}^{3}_{3}$. In Section 5 we shall present the proof of Theorems 1.6-1.8.
The above classification results in $\mathbb{R}_{+}^{n+1}$ enable us to classify nonnegative solutions to biharmonic boundary value problems on $\mathbb{B}^{n+1}$ with one boundary point singularity.
@@
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\end{cases}$ (1.9)
where $i,j\in\{1,2,3\}$ with $i<j$. Assume that there exist $\alpha\in(0,\frac{n(n-3)}{n+3})$ and $C\in\mathbb{R}_{+}$ such that
@@
A direct consequence of Theorem 1.9 is to prove the uniqueness of positive smooth solutions to a certain type of geometric partial differential equations on $\mathbb{B}^{n+1}$.
@@
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\end{cases}$
where $i<j\in\{1,2,3\}$. Then

### v2

@@ 1 Introduction
For $n\geq 2$ and $n\neq 3$, the biharmonic Green functions for $\mathbb{R}_{+}^{n+1}$ are
$G^{(0,2)}(X,Y)=$ $\Gamma(X-Y)-\Gamma(\bar{X}-Y);$
$G^{(1,3)}(X,Y)=$ $\Gamma(X-Y)+\Gamma(\bar{X}-Y);$
$G^{(0,1)}(X,Y)=$ $\Gamma(X-Y)+\frac{1}{2}\Gamma(|\bar{X}-Y|)\left[(n-3)\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-(n-1)\right];$
$G^{(2,3)}(X,Y)=$ $\Gamma(X-Y)-\frac{1}{2}\Gamma(|\bar{X}-Y|)\left[(n-3)\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-(n-1)\right];$
where $\bar{X}=(x,-t)$ for $X\in\mathbb{R}_{+}^{n+1}$. Moreover, there holds
$0\leq G^{(0,1)}(X,Y)\leq G^{(0,2)}(X,Y)\leq G^{(1,3)}(X,Y)\leq G^{(2,3)}(X,Y),\quad\forall~{}X,Y\in\overline{\mathbb{R}_{+}^{n+1}},X\neq Y,$
with equality if and only if either $X\in\mathbb{R}^{n}$ or $Y\in\mathbb{R}^{n}$.
@@
The representation formulae of Green functions for $\mathbb{R}_{+}^{4}$ are
$G^{(0,1)}(X,Y)=$ $\Gamma(X-Y)-\Gamma(\bar{X}-Y)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-1\right);$
$G^{(2,3)}(X,Y)=$ $\Gamma(X-Y)+\Gamma(\bar{X}-Y)-\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|X-Y|^{2}}{|\bar{X}-Y|^{2}}-1\right)+C;$
$G^{(0,2)}(X,Y)=$ $\Gamma(X-Y)-\Gamma(\bar{X}-Y);$
$G^{(1,3)}(X,Y)=$ $\Gamma(X-Y)+\Gamma(\bar{X}-Y)+C;$
for $X,Y\in\overline{\mathbb{R}_{+}^{4}},X\neq Y$ and $C\in\mathbb{R}$.
Since $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$ are conformally equivalent, we extend to introduce the notions of the biharmonic Poisson kernel $\bar{P}_{k}^{3},k\in\{0,1,2,3\}$ and Green function $\bar{G}^{(i,j)},i,j\in\{0,1,2,3\},i<j,i+j\neq 3$ for $\mathbb{B}^{n+1}$, and derive their explicit formulae as well.
Until now, we have succeeded in obtaining a fourth-order analogue of the classical Poisson kernel and Green function for elliptic equations.
For $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$, we assemble the biharmonic Poisson kernel part in Section 3 and the biharmonic Green function part in Section 4.
To continue, we introduce a ‘geometric bubble’ by
$U_{x_{0},\varepsilon}(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\qquad(x,t)\in\mathbb{R}_{+}^{n+1},$
and its restriction to the boundary by letting $U_{x_{0},\varepsilon}(x)=U_{x_{0},\varepsilon}(x,0)$, where $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$.
For $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, under appropriate decay rate of the initial data, the intrinsic $\mathscr{B}^{3}_{3}$ equation with critical/subcritical growth enables us to determine the explicit formula for $u$ on $\partial\mathbb{R}_{+}^{n+1}$, see Lemma 5.1. With the aid of the basics of $U_{x_{0},\varepsilon}$ and biharmonic Poisson integral, we establish that every nonnegative smooth solution is a sum of a bubble and a polynomial. In sharp contrast to second-order elliptic equations, fix $p_{3}=p_{3}^{\ast}$, a striking feature is that a nontrivial bubble (in general, without geometric meaning) exists even for $p_{1}\neq p_{1}^{\ast}$, and coincides with $U_{x_{0},\varepsilon}$ for $p_{1}=p_{1}^{\ast}$; whereas for $p_{1}<p_{1}^{\ast}$, the solution is trivial.
@@
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{1}^{3}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some $c>\max\{3/p_{3},1/p_{1}\}$. Then we have
  1. (1)
    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{1}>\frac{1}{n-3}$, then either $u=0$ or there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{2}$ such that
@@
    $M_{1}(x,t)=\left[\mathbb{T}_{1}^{3}P_{1}^{3}\ast U_{x_{0},\varepsilon}^{p_{1}}+\mathbb{T}_{3}^{3}P_{3}^{3}\ast U_{x_{0},\varepsilon}^{p_{3}^{\ast}}\right](x,t).$
    In particular, when $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
    $u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{2}t^{2}.$
  2. (2)
@@
Although the BVP $(\Delta^{2};\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$ on $\mathbb{R}_{+}^{n+1}$ is not well-posed, instead we take the derivative $\partial_{t}$ of biharmonic Poisson integral of type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ to obtain the integral equation of $u^{p_{1}}(x,0)$, and the classification theorem by Y. Y. Li [30] and Chen-Li-Ou [12]) gives the explicit formula of $u(x,0)$. Our classification result relies not on $p_{1}$ and $p_{2}$ but on the condition whether $p_{2}/p_{1}$ is critical or subcritical Sobolev exponents.
@@
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0,&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{1}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some $c>2/p_{2}$. Then we have
  1. (1)
    If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$ and $p_{1}>0$, then either $u=0$ or there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{3}$ such that
@@
    $M_{2}(x,t)=\Big{[}P_{0}^{3}\ast U_{x_{0},\varepsilon}^{\frac{p_{1}^{\ast}}{p_{1}}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast U_{x_{0},\varepsilon}^{p_{2}^{\ast}}\Big{]}(x,t).$
    In particular, when $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
    $u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{3}t^{3}.$
  2. (2)
@@
For $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$, the analysis of the polynomial part becomes more challenging, and requires the techniques developed in Theorems 1.4 and 1.5.
@@
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\qquad~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume that $u(x,0)=O(|x|^{-c})$ as $|X|\to\infty$ for some $c>\max\{2/p_{2},3/p_{3}\}$. Then we have
  1. (1)
    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then either $u=0$ or there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{1}$ such that
@@
    $M_{3}(x,t)=\left[P_{0}^{3}\ast U_{x_{0},\varepsilon}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast U_{x_{0},\varepsilon}^{p_{2}}\right](x,t).$
    In particular, when $p_{2}=p_{2}^{\ast}=\frac{n+1}{n-3}$, then
    $u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{1}t.$
  2. (2)
@@
In conclusion, $p_{3}$ is the dominant exponent and the dominant impacts of $\mathscr{B}^{3}_{k}$ are different, precisely, $\mathscr{B}^{3}_{1}\preceq\mathscr{B}^{3}_{2}\preceq\mathscr{B}^{3}_{3}$. Especially, for $p_{i}=p_{i}^{\ast},i\in\{1,2,3\}$, the bubble happens to be the same $U_{x_{0},\varepsilon}$; meanwhile, the corresponding polynomial varies according to various circumstances. Theorems 1.4-1.6 give a complete picture of Liouville-type theorems to the study of the $Q$-flat and constant $T$-curvature problem on compact manifolds with boundary in the future.
In Section 2 we study the nontrivial solutions of a homogeneous biharmonic BVP to (1.6) (i.e., $f_{i}=f_{j}=0$) under suitable decay assumption of the initial data, and also describe asymptotic behavior of two typical singular integrals originating from the biharmonic Poisson integrals for $\mathbb{R}_{+}^{n+1}$. In Section 5 we present the proof of classification Theorems 1.4-1.6.
Theorems 1.4-1.6 enable us to classify nonnegative solutions to the $Q$-flat and constant $T$-curvature equations on $\mathbb{B}^{n+1}$ with one boundary point singularity.
@@
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}U=0&\mathrm{~{}~{}in~{}~{}}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\end{cases}$ (1.10)
where $i,j\in\{1,2,3\},i<j$. Assume that there exist $\alpha\in(0,\frac{n(n-3)}{n+3})$ and $C\in\mathbb{R}_{+}$ such that
@@
We would digress for the moment to comment on the decay assumption of initial data that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ in Theorems 1.4-1.6. For critical exponents $p_{i}^{\ast}$, such assumptions are to ensure that involved biharmonic Poisson integrals are well-defined. Consequently, this leads to $c>\frac{3(n-3)}{n+3}$ in Theorems 1.4 ,1.6 and $c>\frac{2(n-3)}{n+1}$ in Theorem 1.5. As will show in the proof of Theorem 1.7, we indeed have $c=n-3-\alpha$, so (1.11) gives $\alpha<\frac{n(n-3)}{n+1}$ and $\alpha<\frac{(n-1)(n-3)}{n+1}$, respectively. Besides this, if $\mathscr{B}^{3}_{3}(U)=\frac{n-3}{2}\tilde{T}^{3}_{3}U^{\frac{n+3}{n-3}}\in L^{1}(\mathbb{S}^{n})$ for a conformal metric $\tilde{g}=U^{4/(n-3)}|\mathrm{d}\xi|^{2}$, then near $-\mathbf{e}_{n+1}$ we have
which requires $0<\alpha<\frac{n(n-3)}{n+3}$. Intuitively, $\mathscr{B}^{3}_{3}(U)\in L^{1}(\mathbb{S}^{n})$ means $U\in W^{3,1}(\mathbb{S}^{n})\hookrightarrow W^{2,\frac{n}{n-1}}(\mathbb{S}^{n})$. This roughly yields $\mathscr{B}^{3}_{2}(U)\in L^{\frac{n}{n-1}}(\mathbb{S}^{n})$, and also
which requires $\alpha<\frac{(n-1)(n-3)}{n+1}$.
@@
$\begin{cases}\displaystyle~{}~{}~{}\Delta^{2}U=0&\mathrm{~{}~{}in~{}~{}}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n},\end{cases}$
where $i,j\in\{1,2,3\},i<j$. Then